OpenStudy (anonymous):

Simplify the radical expression. √14q 2√4q A.3√56q^2 B. 4q√14 C. 2√56q^2 D.4√14q^2

OpenStudy (anonymous):

alright, so pretty much, we can begin by simplifying off the 2√(4q) in the problem everything is being multiplied and 2√(4q)=2*2√a since the square root of 4 is 2. which is why I have another two out there. now, there is some rule in which you can multiply fraction if they have the same root or something like that. So like √5 * √3 = √15 in this case √(14q)*4√q=4√(14q^2) and since we have a squared q inside a square root the q comes out and we get 4q√14

OpenStudy (anonymous):

which is B, understand what I did?

OpenStudy (anonymous):

git it , thanks .

OpenStudy (anonymous):

glad to help